OpenStudy (anonymous):

Solve 2 – 3x = x – 10 x = –3 x = –2 x = 2 x = 3

hartnn (hartnn):

Add 3x to both sides, what u get ?

OpenStudy (anonymous):

x = -2

OpenStudy (hba):

\[2-3x=x-10,-3x-x=-10+2,-4x=-8,x=\frac{ -8 }{ -4 }\]

hartnn (hartnn):

and x is not -2

OpenStudy (hba):

\[x=2\]

hartnn (hartnn):

its not 2 either :P

OpenStudy (anonymous):